OpenStudy (anonymous):

Find the equations of the tangent line and normal line to the curve at the given point. y=x^4+2e^x , (0,2)

OpenStudy (anonymous):

\[ y'= 4 x^3 + 2e ^x\\ y'(0)=2 \]

OpenStudy (anonymous):

Tangent \[ y-2 = 2(x-0)\\ y =2 x + 2 \]

OpenStudy (anonymous):

Normal \[ y -2 =-\frac 1 2 (x-0)\\ y =\frac 1 2 x + 2 \]

OpenStudy (anonymous):

Ok. Now i see it. thx much